The reaction between potassium chlorate and red phosphorus takes place when you strike a match on a matchbox. This happens in the reaction

KClO3 + P4 = P4O10 + KCl

If there are 0.895 g of KClO3 and 0.995 of P4
a. Which is the limiting reactant
b. How much KCl is produced in the reaction?
1
Expert's answer
2021-09-20T02:25:20-0400

3 P4 + 10 KClO3 → 3 P4O10 + 10 KCl

P4 is a reducing agent, KClO3 is an oxidizing agent.


phosphorus is the limiting reactant


Moles of "KClO_3=\frac{0.895}{122.55}=0.0073moles"

Moles of "P_4=\frac{0.995}{124}=0.0080moles"


Moles of "KCl= \frac{10}{3}×0.0080=0.027moles"